### v1

@@ A.2.2 Some technical lemmas
Next we control $\|H_{jk}(\mathbf{x})\nabla_{k}u(\mathbf{x})\|$. Note $\mathbf{D}^{(v)}_{\alpha,i}=0$ if $\alpha$ does not live in block $i$ or a vertical neighbor of block $i$. Similarly for $\mathbf{D}^{(h)}_{\alpha,i}$. Therefore, $H_{ij}(\mathbf{x})=0$ if $i$ and $j$ are not neighbors. For vertical neighbors $i,j$, there are exactly $2m$ boundary indices $\alpha$ s.t. $\left(\mathbf{D}^{(v)}_{\alpha,i}\right)^{\rm T}\mathbf{D}^{(v)}_{\alpha,j}\neq 0$, and $\left\|\left(\mathbf{D}^{(v)}_{\alpha,i}\right)^{\rm T}\mathbf{D}^{(v)}_{\alpha,j}\right\|\leq 2$. Moreover,
$\left\|\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})\mathbf{D}_{\alpha,i}^{(v)}-(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})\mathbf{D}_{\alpha,i}^{(h)}\right)^{\rm T}\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})\mathbf{D}_{\alpha,j}^{(v)}-(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})\mathbf{D}_{\alpha,j}^{(h)}\right)\right\|$
[…]
Therefore, recall Eq. 37, it holds that
(39) $\begin{split}\left\|H_{ij}(\mathbf{x})\right\|\leq~{}&\delta\sum_{\alpha\text{in boundary}}\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}+\varepsilon\right)^{-3/2}\\ &\cdot\left[4\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}\right)+4\varepsilon\right]\\ \leq~{}&4\delta\sum_{\alpha\text{ in boundary}}\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}+\varepsilon\right)^{-1/2}\leq 8m\delta\varepsilon^{-1/2}.\end{split}$
Similar result holds when $i,j$ are horizontal neighbors. Thus for $k\neq j$,
$\|H_{jk}(\mathbf{x})\nabla_{k}u(\mathbf{x})\|\leq 8m\delta\varepsilon^{-1/2}\|\nabla_{k}u(\mathbf{x})\|{\bf{1}}_{j\sim k},$
where $j\sim k$ denotes $j\neq k$ are neighboring blocks. Now substitute the above inequalities into Eq. 38, we have
(40) $\begin{split}&-\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}\nabla_{j}\phi(\mathbf{x}_{i})+\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}\mathcal{L}(\nabla_{j}u)(\mathbf{x})\\ \geq~{}&\lambda M_{jj}\|\nabla_{j}u(\mathbf{x})\|^{2}-\sum_{k\neq j}(\lambda M_{jk}+8m\delta\varepsilon^{-1/2}{\bf{1}}_{j\sim k})\|\nabla_{j}u(\mathbf{x})\|\|\nabla_{k}u(\mathbf{x})\|.\end{split}$
Consider $\mathbf{x}$ where $\|\nabla_{j}u(\mathbf{x})\|$ reaches its maximum, i.e. $\|\nabla_{j}u(\mathbf{x})\|=\left\|\nabla_{j}u\right\|_{L^{\infty}}$. The first order optimality condition reads
$0=\nabla\left(\|\nabla_{j}u(\mathbf{x})\|^{2}\right)=\nabla\nabla_{j}u(\mathbf{x})\cdot\nabla_{j}u(\mathbf{x}),$
and the second order optimality condition reads
$0\geq\Delta\left(\|\nabla_{j}u(\mathbf{x})\|^{2}\right)=~{}$ $2\|\nabla\nabla_{j}u(\mathbf{x})\|_{\rm F}^{2}+2(\nabla_{j}u(\mathbf{x}))^{\rm T}\Delta\nabla_{j}u(\mathbf{x})$
[…]
Therefore,
So that at the maximum point, Eq. 40 reads
$\lambda M_{jj}\|\nabla_{j}u\|_{L^{\infty}}^{2}-$ $\sum_{k\neq j}(\lambda M_{jk}+8m\delta\varepsilon^{-1/2}{\bf{1}}_{j\sim k})\|\nabla_{j}u\|_{L^{\infty}}\|\nabla_{k}u(\mathbf{x})\|$
[…]
Here we use $\|\nabla_{j}\phi(\mathbf{x}_{i})\|\leq\delta_{ij}$, since $\phi$ is $1$-Lipschitz and is only a function of $\mathbf{x}_{i}$. Suppose for the moment that $\|\nabla_{j}u\|_{L^{\infty}}>0$, then it holds
$\delta_{ij}\geq~{}$ $\lambda M_{jj}\|\nabla_{j}u\|_{L^{\infty}}-\sum_{k\neq j}(\lambda M_{jk}+8m\delta\varepsilon^{-1/2}{\bf{1}}_{j\sim k})\|\nabla_{k}u(\mathbf{x})\|$
[…]
Taking summation over those $j\in\mathcal{J}:=\{j\in[b]:\|\nabla_{j}u\|_{L^{\infty}}>0\}$ gives
$1\geq\sum_{j\in\mathcal{J}}\delta_{ij}\geq~{}$ $\sum_{j\in\mathcal{J}}\lambda M_{jj}\|\nabla_{j}u\|_{L^{\infty}}-\sum_{j\in\mathcal{J}}\sum_{k\neq j}(\lambda M_{jk}+8m\delta\varepsilon^{-1/2}{\bf{1}}_{j\sim k})\|\nabla_{k}u\|_{L^{\infty}}$
[…]
@@
So that
$1\geq\frac{\lambda c}{2}\sum_{j\in\mathcal{J}}\|\nabla_{j}u\|_{L^{\infty}}=\frac{\lambda c}{2}\sum_{j=1}^{b}\|\nabla_{j}u\|_{L^{\infty}}.$
The conclusion follows.
@@
$\max_{\alpha}\mathbb{E}_{\mathbf{x}\sim\pi}\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{-1}\leq C_{\pi}.$
@@
$\log\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})=-\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}T^{-1}\mathbf{z}\right\|^{2}-\delta\sum_{\alpha=1}^{d}\left\|\mathbf{D}_{\alpha}(T^{-1}\mathbf{z})\right\|-\log\mu(\mathbf{z}_{\alpha-})+\text{const}.$
Fix $\mathbf{z}_{\alpha-}$ for the moment. Notice $\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}T^{-1}\mathbf{z}\right\|^{2}$ is $L$-smooth for some dimension-independent $L>0$, since only a dimension-independent number of coordinates of $\mathbf{A}T^{-1}\mathbf{z}$ depend on $\mathbf{z}_{\alpha-}$. Therefore, fix any $\mathbf{z}_{\alpha+}^{0}$,
$\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}T^{-1}(\mathbf{z}_{\alpha+},\mathbf{z}_{\alpha-})\right\|^{2}$
[…]
where $\mathbf{v}^{0}$ is the gradient w.r.t. $\mathbf{z}_{\alpha+}$ of $\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}T^{-1}\mathbf{z}\right\|^{2}$ at $\mathbf{z}_{\alpha+}^{0}$. Notice also
$\delta\sum_{\alpha=1}^{d}\left\|\mathbf{D}_{\alpha}(T^{-1}(\mathbf{z}_{\alpha+},\mathbf{z}_{\alpha-}))\right\|\leq\delta\sum_{\alpha=1}^{d}\left\|\mathbf{D}_{\alpha}(T^{-1}(\mathbf{z}_{\alpha+}^{0},\mathbf{z}_{\alpha-}))\right\|+8\delta\left\|\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0}\right\|,$
since change in any one of $\mathbf{z}_{\alpha}^{(v)}$ or $\mathbf{z}_{\alpha}^{(h)}$ affects $4$ finite difference terms in the summation. Combining the above controls, when $\mathbf{z}_{\alpha+}\in B_{1}(\mathbf{z}_{\alpha+}^{0})$ where
$B_{1}(\mathbf{z}_{\alpha+}^{0})=\{\mathbf{z}_{\alpha+}:\left\|\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0}\right\|\leq 1\},$
it holds that
$\log\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})\geq~{}$ $\log\mu(\mathbf{z}_{\alpha+}^{0}|\mathbf{z}_{\alpha-})-\mathbf{v}^{0}\cdot(\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0})$
$-\frac{L}{2}\left\|\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0}\right\|^{2}-8\delta\left\|\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0}\right\|$
[…]
@@
$1=\int\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})\mathrm{d}\mathbf{z}_{\alpha+}\geq\int_{B_{1}(\mathbf{z}_{\alpha+}^{0})}\exp\left(\log\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})\right)\mathrm{d}\mathbf{z}_{\alpha+}$
@@
[…]
where we use the Jensen’s inequality in the last but two line, and use the symmetry of $B_{1}(\mathbf{z}_{\alpha-}^{0})$ in the last equality. Therefore, we prove that
@@
This holds for arbitrary $\mathbf{z}_{\alpha+}^{0}$, so that
$\sup_{\mathbf{z}_{\alpha+}\in\mathbb{R}^{2}}\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})\leq|B_{1}|^{-1}\exp\left(\frac{L}{2}+8\delta\right).$
This implies that the marginal distribution of $\mu(\mathbf{z}_{\alpha+})$ satisfies
@@
$\mathbb{E}_{\mathbf{x}\sim\pi}\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{-1}=~{}$ $\int_{\mathbb{R}^{2}}\frac{\mu(\mathbf{z}_{\alpha+})}{\left\|\mathbf{z}_{\alpha+}\right\|}\mathrm{d}\mathbf{z}_{\alpha}^{(v)}\mathrm{d}\mathbf{z}_{\alpha}^{(h)}$
[…]
@@
A.3 dimension-independent MLwG properties
We will prove that the MLwG algorithm has dimension-independent acceptance rate and convergence rate. The proof is based on the Proposition 3.3 and Theorem 3.6. in [32].
A.3.1 Proof of Proposition 4.1
We need to verify the sparse conditional structure in Assumption 3.1 and the boundedness assumptions in Assumption 3.2 in [32]. The former is trivial, since
$\nabla^{2}_{ij}\log\pi_{\varepsilon}(\mathbf{x})=-(\mathbf{A}^{\rm T}\mathbf{A})(i,j)-\nabla_{ij}^{2}\varphi_{\varepsilon}(\mathbf{x}),$
which is zero if $i$ and $j$ are not neighbors. So that $\max_{j}|\mathcal{I}_{j}|\leq 4$, i.e. one can take the sparsity constant $s=4$.
Next we verify the boundedness condition. (1) Note the vector field is
$\mathbf{v}_{i}(\mathbf{x}):=\nabla_{i}\log\pi_{\varepsilon}(\mathbf{x})=-\mathbf{A}_{\cdot,i}^{\rm T}(\mathbf{A}\mathbf{x}-\mathbf{y})-\nabla_{i}\varphi_{\varepsilon}(\mathbf{x}).$
For the first term, notice $\mathbf{A}_{ji}=0$ if $i$ and $j$ are not neighbors, so that
$\mathbf{A}_{\cdot,i}^{\rm T}(\mathbf{A}\mathbf{x}-\mathbf{y})=\sum_{j\sim i}\mathbf{A}_{ji}^{\rm T}[\mathbf{A}\mathbf{x}-\mathbf{y}]_{j}.$
Denote $R:=R(\mathbf{x})=\max_{j}\|[\mathbf{A}\mathbf{x}-\mathbf{y}]_{j}\|$, then it holds that
$\|\mathbf{A}_{\cdot,i}^{\rm T}(\mathbf{A}\mathbf{x}-\mathbf{y})\|\leq\sum_{j\sim i}\left\|\mathbf{A}_{ji}\right\|\|[\mathbf{A}\mathbf{x}-\mathbf{y}]_{j}\|\leq 4RC_{\mathbf{A}},$
since #$\#\{j:j\sim i\}\leq 4$. For the second term in $\mathbf{v}_{i}(\mathbf{x})$, direct computation shows
$\|\nabla_{i}\varphi_{\varepsilon}(\mathbf{x})\|=~{}$ $\delta\left\|\sum_{\alpha=1}^{d}\frac{(\mathbf{D}^{(h)}_{\alpha}\mathbf{x})\left(\mathbf{D}^{(v)}_{\alpha,i}\right)^{\rm T}+[\mathbf{D}^{(v)}_{\alpha}\mathbf{x}]\left(\mathbf{D}^{(h)}_{\alpha,i}\right)^{\rm T}}{\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}^{(h)}_{\alpha}\mathbf{x})^{2}+\varepsilon\right)^{1/2}}\right\|$
[…]
where we use #$\#\{\alpha:\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|\neq 0\}\leq(m+2)^{2}$, since $\mathbf{D}_{\alpha,j}^{(v)}=\mathbf{D}_{\alpha,j}^{(h)}=0$ if $\alpha$ does not live in block $j$ or a neighbor of block $j$, so that there are at most $(m+2)^{2}$ indices $\alpha$ s.t. $\mathbf{D}_{\alpha}\mathbf{x}$ is nonzero. Therefore, we showed that
$\|\mathbf{v}_{i}(\mathbf{x})\|\leq 4rC_{\mathbf{A}}+2\delta(m+2)^{2}=:M_{\mathbf{v}}.$
(2) For the boundedness of the Hessian, note
$\nabla_{j}\mathbf{v}_{i}(\mathbf{x})=-(\mathbf{A}^{\rm T}\mathbf{A})_{ij}-\nabla_{ij}^{2}\varphi_{\varepsilon}(\mathbf{x}).$
The first term can be bounded by
$\left\|(\mathbf{A}^{\rm T}\mathbf{A})_{ij}\right\|\leq\sum_{k:k\sim i,k\sim j}\left\|\mathbf{A}_{ki}\right\|\left\|\mathbf{A}_{kj}\right\|\leq 4C_{\mathbf{A}}^{2}.$
By Eq. 39, it holds that for
$\left\|\nabla_{ij}^{2}\varphi_{\varepsilon}(\mathbf{x})\right\|\leq 8m\delta\varepsilon^{-1/2}.$
Therefore,
$\left\|\nabla_{j}\mathbf{v}_{i}(\mathbf{x})\right\|\leq 4C_{\mathbf{A}}^{2}+8\delta m\varepsilon^{-1/2}=:H_{\mathbf{v}}.$
(3) For the Lipschitz condition for $\nabla_{j}\mathbf{v}_{i}(\mathbf{x})$, note (see Eq. 37)
$\nabla_{j}\mathbf{v}_{i}(\mathbf{x})-\nabla_{j}\mathbf{v}_{i}(\mathbf{z})=\nabla_{ij}^{2}\varphi_{\varepsilon}(\mathbf{z})-\nabla_{ij}^{2}\varphi_{\varepsilon}(\mathbf{x})$
[…]
where
$\Delta_{1}=g^{(h)}(\mathbf{x})-g^{(h)}(\mathbf{z}),\quad\Delta_{2}=g^{(v)}(\mathbf{x})-g^{(v)}(\mathbf{z}),$
$g^{(h)}(\mathbf{x})=\frac{(\mathbf{D}^{(h)}_{\alpha}\mathbf{x})^{2}+\varepsilon}{\left(\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{2}+\varepsilon\right)^{3/2}},\quad g^{(v)}(\mathbf{x})=\frac{(\mathbf{D}^{(v)}_{\alpha}\mathbf{x})^{2}+\varepsilon}{\left(\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{2}+\varepsilon\right)^{3/2}}.$
Taking derivative, one gets
$\left(\nabla g^{(h)}(\mathbf{x})\right)^{\rm T}=\frac{(\mathbf{D}^{(h)}_{\alpha}\mathbf{x})\mathbf{D}^{(h)}_{\alpha}}{\left(\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{2}+\varepsilon\right)^{3/2}}-\frac{3(\mathbf{D}^{(h)}_{\alpha}\mathbf{x})^{2}\left((\mathbf{D}^{(h)}_{\alpha}\mathbf{x})\mathbf{D}^{(h)}_{\alpha}+(\mathbf{D}^{(v)}_{\alpha}\mathbf{x})\mathbf{D}^{(v)}_{\alpha}\right)}{\left(\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{2}+\varepsilon\right)^{5/2}}.$
Therefore,
$\|\nabla g^{(h)}(\mathbf{x})\|\leq~{}$ $\frac{|\mathbf{D}^{(h)}_{\alpha}\mathbf{x}|\|\mathbf{D}^{(h)}_{\alpha}\|}{\left(\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{2}+\varepsilon\right)^{3/2}}+\frac{3(\mathbf{D}^{(h)}_{\alpha}\mathbf{x})^{2}\left(|\mathbf{D}^{(h)}_{\alpha}\mathbf{x}|\|\mathbf{D}^{(h)}_{\alpha}\|+|\mathbf{D}^{(v)}_{\alpha}\mathbf{x}|\|\mathbf{D}^{(v)}_{\alpha}\|\right)}{\left(\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|^{2}+\varepsilon\right)^{5/2}}$
[…]
Similarly, one can show that $\|\nabla g^{(v)}(\mathbf{x})\|\leq 7\sqrt{2}\varepsilon^{-1}$. Therefore,
$|\Delta_{1}|,|\Delta_{2}|\leq 7\sqrt{2}\varepsilon^{-1}\|\mathbf{x}-\mathbf{z}\|.$
So that it holds
$\left\|\nabla_{j}\mathbf{v}_{i}(\mathbf{x})-\nabla_{j}\mathbf{v}_{i}(\mathbf{z})\right\|$
[…]
This implies we can take the Lipschitz constant $L_{\mathbf{v}}=28\delta(m+2)^{2}\varepsilon^{-1}$.
Now we have verified the sparse conditional structure (Assumption 3.1 in [32]) and the boundedness conditions (Assumption 3.2 in [32]). Then the conclusion follows from Proposition 3.3 in [32], where it is proved of the existence of a dimension-independent constant $M$ s.t. the acceptance rate is bounded below by
$\mathbb{E}\alpha^{k,i}\geq 1-M\sqrt{\tau}.$
A.3.2 Proof of Proposition 4.3
It suffices to verify the blockwise log concavity condition (see Assumption 3.5 in [32]). In Eq. 39, we show that
$\|\nabla_{ij}^{2}\varphi_{\varepsilon}(\mathbf{x})\|\leq 8m\delta\varepsilon^{-1/2}{\bf{1}}_{i\sim j}.$
Since $\mathbf{A}^{\rm T}\mathbf{A}$ is $c$-diagonal block dominant, there exists $M\in\mathbb{R}^{m\times m}$ s.t.
$(\mathbf{A}^{\rm T}\mathbf{A})_{ii}\succeq M_{ii}I,\quad\left\|(\mathbf{A}^{\rm T}\mathbf{A})_{ij}\right\|\leq M_{ij},\quad\sum_{j\neq i}M_{ij}+c\leq M_{ii}.$
Now define $M^{\prime}\in\mathbb{R}^{m\times m}$ s.t.
$M_{ij}^{\prime}:=-\lambda M_{ii}{\bf{1}}_{i=j}+\lambda M_{ij}{\bf{1}}_{i\neq j}+8m\delta\varepsilon^{-1/2}{\bf{1}}_{i\sim j}.$
Notice $\nabla^{2}\log\pi_{\varepsilon}(\mathbf{x})=-\lambda\mathbf{A}^{\rm T}\mathbf{A}-\nabla^{2}\varphi_{\varepsilon}(\mathbf{x})$, then it can directly verified that
$\nabla^{2}\log\pi_{\varepsilon}(\mathbf{x})\preceq M_{ii}^{\prime}I,\quad\|\nabla^{2}\log\pi_{\varepsilon}(\mathbf{x})\|\leq M_{ij}^{\prime},$
It suffices to show that $M^{\prime}$ is negative definite. Notice $-M^{\prime}$ is $\lambda c/2$-diagonal dominant, see Eq. 41. Then a direct application of the Gerŝgorin discs theorem [32] implies that $M^{\prime}$ is negative definite. Since the theorem says that the eigenvalues of $M^{\prime}$ lies in union of the disks
$G(M^{\prime}):=\bigcup_{i=1}^{m}\{\rho\in\mathbb{R}:|\rho-M_{ii}^{\prime}|\leq\sum_{j\neq i}|M_{ij}^{\prime}|\}.$
Here we take $\rho\in\mathbb{R}$ since $M^{\prime}$ is symmetric. Now notice for all $i$,
$\rho\leq M_{ii}^{\prime}+\sum_{j\neq i}|M_{ij}^{\prime}|\leq-\lambda c/2.$
So that $G(M^{\prime})\subset(-\infty,-\lambda c/2]$. This verifies the blockwise log-concavity condition (Assumption 3.5 in [32]). Note we have verified the Assumption 3.1 and 3.2 in the proof of Proposition 4.1 (see Section A.3.1). So that one can apply Theorem 3.6 in [32] to obtain the conclusion.

### v2

@@ A.1.2 Some technical lemmas
Here we use $\|\nabla_{j}\phi(\mathbf{x}_{i})\|_{2}\leq\delta_{ij}$, since $\phi$ is $1$-Lipschitz and is only a function of $\mathbf{x}_{i}$. So if $\|\nabla_{j}u\|_{L^{\infty}}>0$, it holds that
$\delta_{ij}\geq\lambda M_{jj}\|\nabla_{j}u\|_{L^{\infty}}-\sum_{k\neq j}(\lambda M_{jk}+8m\delta\varepsilon^{-1/2}{\bf{1}}_{j\sim k})\|\nabla_{k}u\|_{L^{\infty}}.$
Taking summation over $j\in\mathcal{J}:=\{j\in[b]:\|\nabla_{j}u\|_{L^{\infty}}>0\}$ gives
(40) $1\geq\sum_{j\in\mathcal{J}}\delta_{ij}\geq\sum_{j\in\mathcal{J}}\|\nabla_{j}u\|_{L^{\infty}}\left[\lambda M_{jj}-\sum_{k\neq j}(\lambda M_{kj}+8m\delta\varepsilon^{-1/2}{\bf{1}}_{j\sim k})\right].$
@@
Together with (40), we obtain (34).
@@
$\max_{\alpha}\mathbb{E}_{\mathbf{x}\sim\pi}\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{-1}\leq C_{\pi}.$
@@
$\log\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})=-\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}T^{-1}\mathbf{z}\right\|_{2}^{2}-\delta\sum_{\alpha=1}^{d}\left\|\mathbf{D}_{\alpha}(T^{-1}\mathbf{z})\right\|_{2}-\log\mu(\mathbf{z}_{\alpha-})+\text{const}.$
Fix $\mathbf{z}_{\alpha-}$ for the moment. Notice $\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}T^{-1}\mathbf{z}\right\|_{2}^{2}$ is $L$-smooth for some dimension-independent $L>0$, since only a dimension-independent number of coordinates of $\mathbf{A}T^{-1}\mathbf{z}$ depend on $\mathbf{z}_{\alpha-}$. Therefore, fix any $\mathbf{z}_{\alpha+}^{0}$,
$\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}T^{-1}(\mathbf{z}_{\alpha+},\mathbf{z}_{\alpha-})\right\|_{2}^{2}-\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}T^{-1}(\mathbf{z}_{\alpha+}^{0},\mathbf{z}_{\alpha-})\right\|_{2}^{2}-\mathbf{v}^{0}\cdot(\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0})$
$\leq\frac{L}{2}\left\|\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0}\right\|_{2}^{2}$ ,
where $\mathbf{v}^{0}$ is the gradient w.r.t. $\mathbf{z}_{\alpha+}$ of $\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}T^{-1}\mathbf{z}\right\|_{2}^{2}$ at $\mathbf{z}_{\alpha+}^{0}$. Notice also
$\delta\sum_{\alpha=1}^{d}\left\|\mathbf{D}_{\alpha}(T^{-1}(\mathbf{z}_{\alpha+},\mathbf{z}_{\alpha-}))\right\|_{2}\leq\delta\sum_{\alpha=1}^{d}\left\|\mathbf{D}_{\alpha}(T^{-1}(\mathbf{z}_{\alpha+}^{0},\mathbf{z}_{\alpha-}))\right\|_{2}+8\delta\left\|\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0}\right\|_{2},$
since change in any one of $\mathbf{z}_{\alpha}^{(v)}$ or $\mathbf{z}_{\alpha}^{(h)}$ affects $4$ finite difference terms in the summation. Combining the above controls, when $\mathbf{z}_{\alpha+}\in B_{1}(\mathbf{z}_{\alpha+}^{0}):=\{\mathbf{z}_{\alpha+}:\left\|\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0}\right\|_{2}\leq 1\}$, it holds that
$\log\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})-$ $\log\mu(\mathbf{z}_{\alpha+}^{0}|\mathbf{z}_{\alpha-})+\mathbf{v}^{0}\cdot(\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0})$
$\geq$ $-\frac{L}{2}\left\|\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0}\right\|_{2}^{2}-8\delta\left\|\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0}\right\|_{2}\geq-\frac{L}{2}-8\delta.$
@@
$1=~{}$ $\int\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})\mathrm{d}\mathbf{z}_{\alpha+}\geq\int_{B_{1}(\mathbf{z}_{\alpha+}^{0})}\exp\left(\log\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})\right)\mathrm{d}\mathbf{z}_{\alpha+}$
@@
where we use the Jensen’s inequality and the symmetry of $B_{1}(\mathbf{z}_{\alpha-}^{0})$. Therefore,
@@
This holds for arbitrary $\mathbf{z}_{\alpha+}^{0}$, so that the marginal distribution of $\mu(\mathbf{z}_{\alpha+})$ satisfies
@@
$\mathbb{E}_{\mathbf{x}\sim\pi}\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{-1}=~{}$ $\int_{\mathbb{R}^{2}}\frac{\mu(\mathbf{z}_{\alpha+})}{\left\|\mathbf{z}_{\alpha+}\right\|_{2}}\mathrm{d}\mathbf{z}_{\alpha}^{(v)}\mathrm{d}\mathbf{z}_{\alpha}^{(h)}\leq 1+\int_{\left\|\mathbf{z}_{\alpha+}\right\|_{2}\leq 1}\frac{C^{\prime}}{\left\|\mathbf{z}_{\alpha+}\right\|_{2}}\mathrm{d}\mathbf{z}_{\alpha}^{(v)}\mathrm{d}\mathbf{z}_{\alpha}^{(h)}$
@@
A.2 Proof of Lemma 5.1